OpenStudy (anonymous):

If m/n = 5/6, then what is 3m + 2n?

OpenStudy (anonymous):

m/n = 5/6 m = 5 n = 6 3m + 2n 3(5) + 2(6) 15 + 12 27

OpenStudy (anonymous):

rearrange the m/n = 5/6 in terms of the variable you like